A water tank is of height H and is comleted filled ,there is hole of area of cross section A at the base . how much water comes from it in one second

Grade:12th pass

1 Answers

Vikas TU
7 years ago
At the bottom the speed would be => root(2gh) => root(2*9.8*H) = > root(19.6H)
And in one second the water that would be come out is:
x = v*1 = > root(19.6H) m.